# A mesoscopic model for the rheology of dilute and semidilute solutions of wormlike micelles

### v1

@@ II.3.1 Fusion and scission rules
where $\theta$ is the angle between the adjacent springs at junction points (for instance, $\theta_{2}$ in Fig. 2B). Note that the function, $f(r)=({1}/{2})\left[\cos\left(\alpha\,{\left({r}/{\sigma}\right)}^{2}+\beta\right)-1\right]$, which is the prefactor to the sticker energy term in Eq. (19), has the values $f(r)=-1$ for $r=2^{1/6}\sigma$ and $f(r)=0$ for $r=r_{c}$. A pseudo-random number is drawn from a uniform distribution between 0 and 1, and if the random number is less than $\exp\left(-\Delta E/k_{\text{B}}T\right)$, the change of the state is accepted and the bond is formed. In an update sweep, each existing bond attempts to break in this manner, and then bond formations are attempted. Further, if two stickers are within the cutoff distance but at least one of them already has as many bonds as its functionality allows, then bond formation for the new pair is not attempted. Here, as mentioned earlier, the functionality of a sticker is set to one, which avoids branching. The use of Eq. (19) to calculate the change in energy ensures that persistent worms are more likely to stick when they are highly aligned and that detailed balance is satisfied, meaning all bond states are reversible and have complementary probabilities. As noted earlier, the breakage and reforming of wormlike micelles is queried once before each Brownian dynamics timestep.
@@ II.3.3 Simulation Details
[…]
@@ II.4 Property definitions
[…]
@@ III.1 Validation of bending potential implementation
Two properties are compared across the two systems. The first property is the distribution $\psi(F^{\text{(b)}}_{\mu})$ of the magnitude of the bending force $F^{\text{(b)}}_{\mu}=|\bm{F}^{\text{(b)}}_{\mu}|$ acting on each of the individual beads within the linear wormlike micelle, and the corresponding beads on the semi-flexible homopolymer chain. As displayed in Fig. 5 (a) to (d), for two different values of the bending stiffness represented by the ratios $\ell_{\text{pw}}/l_{p}=0.01$ and $\ell_{\text{pw}}/l_{p}=1$, the force distributions on the respective beads are nearly identical in both the systems. It is worth noting that for each system, the force distributions for beads 1 and 5, and beads 2 and 4 lie on top of each other, as expected from symmetry considerations. All the remaining bead pairs in the two systems, which are not displayed for the sake of brevity, behave identically for both values of the bending stiffness.
@@ III.2 The overlap concentration
Following the procedure outlined in Sec. II.5, the dependence of $L^{*}_{\text{c}}$ and $\bar{L}$ on $c^{\text{eff}}$, for various values of sticker energy $\epsilon_{st}$, was computed by varying $c/c^{*}_{\text{pw}}$ for systems with and without rings, and plotted together, as shown in Fig. 6 (a). As discussed in Sec. II.5, the intersection point of the two functions $L^{*}_{\text{c}}(c^{\text{eff}})$ and $\bar{L}(c^{\text{eff}})$, for any sticker energy, determines the pair $(L^{*},c^{*})$ for that value of $\epsilon_{st}$. While, as expected, $L^{*}_{\text{c}}$ is a monotonically decreasing function of $c^{\text{eff}}$, the dependence of $\bar{L}$ on $c^{\text{eff}}$ clearly displays two distinct power law regimes that lie on either side of the intersection point, which corresponds to the overlap concentration, with slopes represented by $\alpha_{\text{d}}=0.46$ and $\alpha_{\text{s}}=0.6$. A more thorough discussion of this interesting observation is postponed to Sec. III.3.2, where it interpreted in the light of analytical mean-field theories. Notably, in systems in which both linear chains and rings are present, the presence of rings makes no difference to the estimation of $c^{*}$. The scaled concentration $c^{\text{eff}}/c^{*}$, determined in this manner, for various values of $c/c^{*}_{\text{pw}}$ and $\epsilon_{st}$, is displayed in Table 2, for systems with and without rings.
@@
Table 2: The scaled concentration $c^{\text{eff}}/c^{*}$ determined by the procedure discussed in Sec. II.5 and Fig. 4, for various values of $c/c^{*}_{\text{pw}}$ and $\epsilon_{st}$, for systems with (WR) and without rings (NR).
$c^{\text{eff}}/c^{*}$
$c/c^{*}_{\text{pw}}$ $\epsilon_{st}=4$ $\epsilon_{st}=5$ $\epsilon_{st}=6$ $\epsilon_{st}=7$ $\epsilon_{st}=8$ $\epsilon_{st}=9$ $\epsilon_{st}=10$
NR WR NR WR NR WR NR WR NR WR NR WR NR WR
0.10 0.17 0.13 0.19 0.13 0.23 0.12 0.27 0.10 0.32 0.10 0.39 0.10 0.49 0.10
0.20 0.32 0.26 0.36 0.27 0.43 0.29 0.51 0.31 0.63 0.35 0.79 0.39 0.97 0.47
0.30 0.46 0.39 0.53 0.43 0.62 0.47 0.75 0.53 0.93 0.62 1.16 0.73 1.46 0.90
0.40 0.60 0.53 0.69 0.59 0.83 0.67 1.00 0.76 1.24 0.90 1.55 1.10 1.94 1.35
0.50 0.74 0.66 0.86 0.74 1.02 0.86 1.24 0.99 1.53 1.20 1.93 1.46 2.42 1.83
0.60 0.88 0.79 1.02 0.89 1.22 1.04 1.48 1.22 1.84 1.51 2.31 1.81 2.90 2.29
0.70 1.01 0.92 1.18 1.05 1.41 1.23 1.71 1.45 2.13 1.79 2.69 2.16 3.38 2.77
1.00 1.42 1.32 1.66 1.52 1.99 1.80 2.43 2.15 3.04 2.67 3.83 3.26 5.39 4.19
@@
In this work, the procedure described in Sec. II.5 is used to determine the overlap concentration (displayed in Fig. 6 (a) and Table 2). As will be demonstrated in subsequent sections, the use of this value $c^{*}$ to separate the dilute and semidilute concentration regimes leads to results that are completely consistent with the predictions by analytical mean-field theories.
@@ III.3.1 Analytical predictions of static properties
Figure 8: Probability distribution of lengths of linear wormlike micelles for systems with (WR) and without (NR) rings, at various values of the sticker energy $\epsilon_{st}$, in (a) dilute and (b) semidilute solutions. The concentration is reported in terms of $c/c^{*}_{\text{pw}}$, which is held constant in the two regimes. Each set of symbols at a fixed value of $\epsilon_{st}$ corresponds to a unique value of scaled concentration $c^{\text{eff}}/c^{*}$, which is related to $c/c^{*}_{\text{pw}}$ and $\epsilon_{st}$ as listed in Table 2.
@@ III.3.2 Length distributions and the mean length of linear wormlike micelles
Figure 9: The dependence of the ratio of the mean length $\bar{L}$ to the overlap length $L^{*}$ on the scaled concentration $c^{\text{eff}}/c^{*}$ for systems with and without rings, at various values of the sticker energy $\epsilon_{st}$. The collapse of data, and the values of the power law exponents in the two regimes agree with the predictions of mean-field theory.
@@
According to mean-field theory, when the mean length $\bar{L}$ is scaled by the overlap length $L^{*}$ and plotted as a function of $c^{\text{eff}}/c^{*}$, data for different sticker energies is expected to collapse on to a master plot, as expressed in Eq. (37), with the exponent $\alpha$ assuming values of $0.46$ and $0.6$ in the dilute and semidilute regimes, respectively. Fig. 9 is a plot of simulation results in terms of these variables, and the collapse of data for different sticker energies is clearly visible, with the expected values of the exponents in the two concentration regimes. The scatter in the data at low concentrations and sticker energies is because the mean length of micelles is too short to obey scaling theories. The collapse of data when represented in terms of scaled variables, and the scatter at low values of $c^{\text{eff}}/c^{*}$ and $\epsilon_{st}$, mirrors the identical observations made earlier in their Monte Carlo simulations by Wittmer et al. [72].
The overlap concentration $c^{*}$ and the overlap length $L^{*}$ have been computed as described in Sec. II.5, with the results displayed earlier in Fig. 6 for various values of $\epsilon_{st}$. The power law exponents of $0.46$ and $0.6$ for the dependence of $\bar{L}$ on $c^{\text{eff}}$, seen in Fig. 6 (a), can now be understood as a validation of the prediction by mean-field theory (see Eq. (37)). The crossover from the dilute to the semidilute regime in Fig. 9 occurs fairly sharply at $\bar{L}/L^{*}=c^{\text{eff}}/c^{*}=1$, indicating that the procedure for estimating $c^{*}$ accurately captures the demarcation of the dilute and semidilute regimes. The data collapse for systems with and without rings also indicates that the presence of rings makes no difference to the static properties of linear wormlike micelles, in line with a similar observation made earlier by Wittmer et al. [73].
@@
Since the value of the exponent $\alpha$ in the two concentration regimes has been determined, the amplitudes $A_{d}$ and $A_{s}$, and exponents $\delta_{d}$ and $\delta_{s}$ in Eq. (37), can be determined by plotting $\bar{L}/\left(c^{\text{eff}}\right)^{\alpha}$ versus $\epsilon_{st}$ in the two regimes, as displayed in Fig. 10 (a) and (b). The values of $\delta_{d}=0.46$ and $\delta_{s}=0.5$, determined from data at sufficiently high $\epsilon_{st}$ (where scaling laws are obeyed), are in complete agreement with the predictions of mean-field theory. The expectation that both $c^{*}$ and $L^{*}$ are exponential functions of $\epsilon_{st}$ is borne out by Fig. 10 (c) and (d). The exponents $\varphi$ and $\kappa$ in Eq. (III.3.1) and the prefactors $P$ and $Q$, can be determined by fitting the data in these figures. The value of $\varphi$ is found to be $3.70\pm 0.20$ and that of $\kappa$ is found to be $3.00\pm 0.02$. These values are reasonably close to the values of $3.5$ and $2.8$ expected from mean-field theory. The prefactors $A_{d}=15.74\pm 0.10$ and $A_{s}=22.11\pm 0.50$, and the amplitudes $P=0.09\pm 0.02$ and $Q=5.2\pm 0.2$ can be seen to satisfy the relations given in Eq. (III.3.1). Once the mean length is determined, it is possible to plot the probability distribution $p(x)$ as a function of the scaled length $x$ at various concentrations. Fig. 11 (a) to (c) display $p(x)$ as a function of $x$ in the dilute regime at $c^{\text{eff}}/c^{*}=0.3$, at the overlap concentration, $c^{\text{eff}}/c^{*}=1$, and in the semidilute regime at $c^{\text{eff}}/c^{*}=2$, respectively. In all the cases, the sticker energy is held fixed at $\epsilon_{st}=7$, and systems with and without rings have been computed. The simulation data is fitted with the expression for the Schulz-Zimm distribution, given in Eq. (36). Note that when $\gamma=1$, the Schulz-Zimm distribution reduces to a purely exponential distribution. It is possible, consequently to fit simulation data for various values of $c^{\text{eff}}/c^{*}$ by using $\gamma$ as a fitting parameter. The value of $\gamma$ that fits the data at any concentration is denoted here as $\gamma_{\text{eff}}$. It is clear from Fig. 11 (a) that the Schulz-Zimm distribution is obeyed in the dilute solution, with the exponent $\gamma_{\text{eff}}=1.162\pm 0.002$, which is close to the value expected from mean-field theory. In a semidilute solution at $c^{\text{eff}}/c^{*}=2$, the exponent $\gamma_{\text{eff}}=1.00\pm 0.02$, suggesting that the probability distribution is a pure exponential function, as predicted by mean-field theory. At intermediate concentrations, the value of $\gamma_{\text{eff}}$ lies between $1.162$ and $1$, and decreases with increasing concentration. Fig. 11 (b) indicates that $\gamma_{\text{eff}}=1.10\pm 0.02$ at $c^{\text{eff}}/c^{*}=1$. Similar behaviour was observed for other values of the sticker energy, and Fig. 11 (d), which is a plot of $\gamma_{\text{eff}}$ versus $c^{\text{eff}}/c^{*}$, shows that the crossover from the dilute to the semidilute concentration regime occurs roughly between $c^{\text{eff}}/c^{*}=1$ and $2$ in all cases.
Figure 12: (a) Independence of the probability distribution of scaled lengths $p(x)$ from the choice of the model for the persistent worm, whether it is a trumbbell (T) with spring rest lengths $b=2$ or $3$, or a dumbbell (D) with $b=2^{1/6}$, $2$ or $3$. (b) Polydispersity index $I$ as a function of scaled concentration for systems with and without rings, at various values of the sticker energy $\epsilon_{st}$. Data is shown only for cases where ${\text{c}}^{\text{eff}}/\text{c}^{*}>0.2$ and $\bar{L}>20$, where the scaling laws are obeyed.
@@
Figure 13: Probability distribution of lengths of ring wormlike micelles, at various values of the sticker energy $\epsilon_{st}$, in (a) dilute and (b) semidilute solutions. The concentration is reported in terms of $c/c^{*}_{\text{pw}}$, which is held constant in the two regimes. Each set of symbols at a fixed value of $\epsilon_{st}$ corresponds to a unique value of scaled concentration $c^{\text{eff}}/c^{*}$, which is related to $c/c^{*}_{\text{pw}}$ and $\epsilon_{st}$ as listed in Table 2.

### v2

@@ II.3.1 Fusion and scission rules
where $\theta$ is the angle between the adjacent springs at junction points (for instance, $\theta_{2}$ in Fig. 3B). Note that the function, $f(r)=({1}/{2})\left[\cos\left(\alpha\,{\left({r}/{\sigma}\right)}^{2}+\beta\right)-1\right]$, which is the prefactor to the sticker energy term in Eq. (19), has the values f(r)=1f(r)=-1 for $r=2^{1/6}\sigma$ and $f(r)=0$ for $r=r_{c}$. A pseudo-random number is drawn from a uniform distribution between 0 and 1, and if the random number is less than $\exp\left(-\Delta E/k_{\text{B}}T\right)$, the change of the state is accepted and the bond is formed. In an update sweep, each existing bond attempts to break in this manner, and then bond formations are attempted. Further, if two stickers are within the cutoff distance but at least one of them already has as many bonds as its functionality allows, then bond formation for the new pair is not attempted. Here, as mentioned earlier, the functionality of a sticker is set to one, which avoids branching. The use of Eq. (19) to calculate the change in energy ensures that persistent worms are more likely to stick when they are highly aligned and that detailed balance is satisfied, meaning all bond states are reversible and have complementary probabilities. As noted earlier, the breakage and reforming of wormlike micelles is queried once before each Brownian dynamics timestep. Simulations have been carried out at various decreasing timesteps to ensure convergence of dynamic property predictions, and reported values are independent of the choice of timestep.
@@ II.3.3 Simulation Details
1 Backbone interaction strength $\text{$\epsilon$}_{bb}$ 0
2 Sticker strength $\text{$\epsilon$}_{st}$ 2 to 10
@@ II.4 Property definitions
and angular brackets $\langle\cdots\rangle$ represent an ensemble average over all stochastic trajectories. The radius of gyration of ring wormlike micelles, $R_{g_{R}}$, is calculated similarly, with the number of effective monomers $N_{\text{L,R}}^{\text{eff}}$ in a ring wormlike micelle of length $L$, taking the place of $N_{\text{L,lin}}^{\text{eff}}$ in the expressions above.
@@ III.1 Validation of bending potential implementation
[…]
@@ III.2 The overlap concentration
Following the procedure outlined in Sec. II.5, the dependence of $L^{*}_{\text{c}}$ and $\bar{L}$ on $c^{\text{eff}}$, for various values of sticker energy $\epsilon_{st}$, was computed by varying $c/c^{*}_{\text{pw}}$ for systems with and without rings, and plotted together, as shown in Fig. 7 (a). As discussed in Sec. II.5, the intersection point of the two functions $L^{*}_{\text{c}}(c^{\text{eff}})$ and $\bar{L}(c^{\text{eff}})$, for any sticker energy, determines the pair $(L^{*},c^{*})$ for that value of $\epsilon_{st}$. While, as expected, $L^{*}_{\text{c}}$ is a monotonically decreasing function of $c^{\text{eff}}$, the dependence of $\bar{L}$ on $c^{\text{eff}}$ clearly displays two distinct power law regimes that lie on either side of the intersection point, which corresponds to the overlap concentration, with slopes represented by $\alpha_{\text{d}}=0.46$ and $\alpha_{\text{s}}=0.6$. A more thorough discussion of this interesting observation is postponed to Sec. III.3.2, where it interpreted in the light of analytical mean-field theories. Notably, in systems in which both linear chains and rings are present, the presence of rings makes no difference to the estimation of $c^{*}$.
The scaled concentration $c^{\text{eff}}/c^{*}$, determined in this manner, for various values of $c/c^{*}_{\text{pw}}$ and $\epsilon_{st}$, is displayed in Figs. 8 (a) and (b), for systems with and without rings. The plots show that $c^{\mathrm{eff}}/c^{*}$ is significantly higher for wormlike micellar solutions containing only linear chains (No Rings) compared to those with blend of linear chains and rings (With Rings), and this difference increases with the sticker energy $\epsilon_{st}$.
Figure 8: Plots of $c^{\mathrm{eff}}/c^{*}$ versus $c/c_{\mathrm{pw}}^{*}$ for various sticker energies comparing systems without rings (NR, filled symbols) and with rings (WR, open symbols). (a) $\epsilon_{st}\leq 6$, (b) $\epsilon_{st}>6$.
@@
In this work, the procedure described in Sec. II.5 is used to determine the overlap concentration (displayed in Fig. 7 (a) and Figs. 8). As will be demonstrated in subsequent sections, the use of this value $c^{*}$ to separate the dilute and semidilute concentration regimes leads to results that are completely consistent with the predictions by analytical mean-field theories.
@@ III.3.1 Analytical predictions of static properties
Figure 10: Probability distribution of lengths of linear wormlike micelles for systems with (WR) and without (NR) rings, at various values of the sticker energy $\epsilon_{st}$, in (a) dilute and (b) semidilute solutions. The concentration is reported in terms of $c/c^{*}_{\text{pw}}$, which is held constant in the two regimes. Each set of symbols at a fixed value of $\epsilon_{st}$ corresponds to a unique value of scaled concentration $c^{\text{eff}}/c^{*}$, which is related to $c/c^{*}_{\text{pw}}$ and $\epsilon_{st}$ as displayed in Fig. 8.
@@ III.3.2 Length distributions and the mean length of linear wormlike micelles
Figure 11: The dependence of the ratio of the mean length $\bar{L}$ to the overlap length $L^{*}$ on the scaled concentration $c^{\text{eff}}/c^{*}$ for systems with (WR) and without rings (NR), at various values of the sticker energy $\epsilon_{st}$. The collapse of data, and the values of the power law exponents in the two regimes agree with the predictions of mean-field theory.
According to mean-field theory, when the mean length $\bar{L}$ is scaled by the overlap length $L^{*}$ and plotted as a function of $c^{\text{eff}}/c^{*}$, data for different sticker energies is expected to collapse on to a master plot, as expressed in Eq. (37), with the exponent $\alpha$ assuming values of $0.46$ and $0.6$ in the dilute and semidilute regimes, respectively. Fig. 11 is a plot of simulation results in terms of these variables, where the data is plotted as $\left(\bar{L}/L^{*}\right)/\left[(c^{\mathrm{eff}}/c^{*})^{0.46}\right]$ versus $c^{\mathrm{eff}}/c^{*}$ to better illustrate the crossover from the dilute to the semidilute regime. The collapse of data for different sticker energies is clearly visible, with the expected values of the exponents in the two concentration regimes. The gradual change in the effective exponent occurs in the concentration range $0.7\lesssim c^{\mathrm{eff}}/c^{*}\lesssim 1$. The scatter in the data at low concentrations and sticker energies is because the mean length of micelles is too short to obey scaling theories. The collapse of data when represented in terms of scaled variables, and the scatter at low values of $c^{\text{eff}}/c^{*}$ and $\epsilon_{st}$, mirrors the identical observations made earlier in their Monte Carlo simulations by Wittmer et al. [75].
@@
The overlap concentration $c^{*}$ and the overlap length $L^{*}$ have been computed as described in Sec. II.5, with the results displayed earlier in Fig. 7 for various values of $\epsilon_{st}$. The power law exponents of $0.46$ and $0.6$ for the dependence of $\bar{L}$ on $c^{\text{eff}}$, seen in Fig. 7 (a), can now be understood as a validation of the prediction by mean-field theory (see Eq. (37)). The data collapse for systems with and without rings also indicates that the presence of rings makes no difference to the static properties of linear wormlike micelles, in line with a similar observation made earlier by Wittmer et al. [76].
@@
Since the value of the exponent $\alpha$ in the two concentration regimes has been determined, the amplitudes $A_{d}$ and $A_{s}$, and exponents $\delta_{d}$ and $\delta_{s}$ in Eq. (37), can be determined by plotting $\bar{L}/\left(c^{\text{eff}}\right)^{\alpha}$ versus $\epsilon_{st}$ in the two regimes, as displayed in Fig. 12 (a) and (b). The values of $\delta_{d}=0.46$ and $\delta_{s}=0.5$, determined from data at sufficiently high $\epsilon_{st}$ (where scaling laws are obeyed), are in complete agreement with the predictions of mean-field theory. The expectation that both $c^{*}$ and $L^{*}$ are exponential functions of $\epsilon_{st}$ is borne out by Fig. 12 (c) and (d). The exponents $\varphi$ and $\kappa$ in Eq. (III.3.1) and the prefactors $P$ and $Q$, can be determined by fitting the data in these figures. The value of $\varphi$ is found to be $3.70\pm 0.20$ and that of $\kappa$ is found to be $3.00\pm 0.02$. These values are reasonably close to the values of $3.5$ and $2.8$ expected from mean-field theory. The prefactors $A_{d}=15.74\pm 0.10$ and $A_{s}=22.11\pm 0.50$, and the amplitudes $P=0.09\pm 0.02$ and $Q=5.2\pm 0.2$ can be seen to satisfy the relations given in Eq. (III.3.1).
Once the mean length is determined, it is possible to plot the probability distribution $p(x)$ as a function of the scaled length $x$ at various concentrations. Fig. 13 (a) to (c) display $p(x)$ as a function of $x$ in the dilute regime at $c^{\text{eff}}/c^{*}=0.3$, at the overlap concentration, $c^{\text{eff}}/c^{*}=1$, and in the semidilute regime at $c^{\text{eff}}/c^{*}=2$, respectively. In all the cases, the sticker energy is held fixed at $\epsilon_{st}=7$, and systems with and without rings have been computed. The simulation data is fitted with the expression for the Schulz-Zimm distribution, given in Eq. (36). Note that when $\gamma=1$, the Schulz-Zimm distribution reduces to a purely exponential distribution. It is possible, consequently to fit simulation data for various values of $c^{\text{eff}}/c^{*}$ by using $\gamma$ as a fitting parameter. The value of $\gamma$ that fits the data at any concentration is denoted here as $\gamma_{\text{eff}}$. It is clear from Fig. 13 (a) that the Schulz-Zimm distribution is obeyed in the dilute solution, with the exponent $\gamma_{\text{eff}}=1.162\pm 0.002$, which is close to the value expected from mean-field theory. In a semidilute solution at $c^{\text{eff}}/c^{*}=2$, the exponent $\gamma_{\text{eff}}=1.00\pm 0.02$, suggesting that the probability distribution is a pure exponential function, as predicted by mean-field theory. At intermediate concentrations, the value of $\gamma_{\text{eff}}$ lies between $1.162$ and $1$, and decreases with increasing concentration. Fig. 13 (b) indicates that $\gamma_{\text{eff}}=1.10\pm 0.02$ at $c^{\text{eff}}/c^{*}=1$. Similar behaviour was observed for other values of the sticker energy, and Fig. 13 (d), which is a plot of $\gamma_{\text{eff}}$ versus $c^{\text{eff}}/c^{*}$, shows that the crossover from the dilute to the semidilute concentration regime occurs roughly between $c^{\text{eff}}/c^{*}=1$ and $2$ in all cases.
@@
Figure 14: (a) Independence of the probability distribution of scaled lengths $p(x)$ from the choice of the model for the persistent worm, whether it is a trumbbell (T) with spring rest lengths $b=2$ or $3$, or a dumbbell (D) with $b=2^{1/6}$, $2$ or 33. (b) Polydispersity index II as a function of scaled concentration for systems with and without rings, at various values of the sticker energy $\epsilon_{st}$. Data is shown only for cases where ${\text{c}}^{\text{eff}}/\text{c}^{*}>0.2$ and $\bar{L}>20$, where the scaling laws are obeyed.
@@
Figure 15: Probability distribution of lengths of ring wormlike micelles, at various values of the sticker energy $\epsilon_{st}$, in (a) dilute and (b) semidilute solutions. The concentration is reported in terms of $c/c^{*}_{\text{pw}}$, which is held constant in the two regimes. Each set of symbols at a fixed value of $\epsilon_{st}$ corresponds to a unique value of scaled concentration $c^{\text{eff}}/c^{*}$, which is related to $c/c^{*}_{\text{pw}}$ and $\epsilon_{st}$ as displayed in Fig. 8.